OpenStudy (anonymous):

please HELPPPPPP!!!!!!??!! DFONT TELL ANSWER, EXPLAIN FIRST!! Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid 135$ . This year, he paid162$. What was the percent increase in the cost of his plane ticket? Show your work

OpenStudy (anonymous):

@cwtan why did you delet what you put??

OpenStudy (istim):

Wouldn't you just divide the initial amount by the second amount?

OpenStudy (cwtan):

I found that I might be wrong

OpenStudy (anonymous):

well do i divide the amounts?

OpenStudy (mertsj):

Why did you post this again? I told you before to find the increase by subtracting. Then divide the increase by the original amount and change the answer to a percent.

OpenStudy (anonymous):

@Mertsj because i didnt understand what you were saying , but now i do ,

OpenStudy (anonymous):

so the answer would be 20% right???

OpenStudy (anonymous):

Ok. First you have to find the amount the price increased. You do that by finding the difference between 162 and 135. Then, you divide this number by 162. Then multiply by 100.

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

135-162=27/162*100=16% ??

OpenStudy (anonymous):

16.67% to be precise, but yes. You find the difference in order to find the amount increased. Then you divide it by the whole in order to get the decimal value. Then, you multiply by 100 to get it from a decimal to a percentage.